PLEASE HELP!! I don't understand how to figure this out. Do I have to find the vertex? Put it in a quadratic equation????
Parker’s Skateboard Shop has determined that when x skateboards are built, the average cost C per skateboard is given by C=x^2-80x+1625 . How many skateboards should the shop build in order to minimize the average cost per skateboard?
Tips: Since # skateboards = x in our situation, find the x-value of the minimum (vertex).

What is the minimum average cost?
Tips: Plug in 2a into cost equation

Answer by scott8148(6628)

the vertex is on the axis of symmetry

x = -b / (2a) = 80 / (2 * 1) = 40

C = 40^2 - 80(40) + 1625